OpenStudy (anonymous):

Jill is 6 years older than her brother ryan.write an equation relating jills age to ryans age . explain what each varible represents.use your equation find ryans age when jill is 14.

Directrix (directrix):

Let J = Jill's age and R = Ryan's age J = R + 6 If J = 14, then 14 = R + 6 14 = R + 6 14 - 6 = R 8 = R